# Follow My Instruction and Spill the Beans: Scalable Data Extraction from Retrieval-Augmented Generation Systems

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) improves pre-trained models by incorporating external knowledge at test time to enable customized adaptation. We study the risk of datastore leakage in Retrieval-In-Context RAG Language Models (LMs). We show that an adversary can exploit LMs’ instruction-following capabilities to easily extract text data verbatim from the datastore of RAG systems built with instruction-tuned LMs via prompt injection. The vulnerability exists for a wide range of modern LMs that span Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and the exploitability exacerbates as the model size scales up. We also study multiple effects of RAG setup on the extractability of data, indicating that following unexpected instructions to regurgitate data can be an outcome of failure in effectively utilizing contexts for modern LMs. Extending our study to production RAG models GPTs, we design an attack that can cause datastore leakage with a 100% success rate on 25 randomly selected customized GPTs with at most 2 queries, and we extract text data verbatim at a rate of 41% from a book of 77,000 words and 3% from a corpus of 1,569,000 words by prompting the GPTs with only 100 queries generated by themselves. Code and data are avaialbe at https://github.com/zhentingqi/rag-privacy.
@@ 1 Introduction
In particular, the inherent propensity of large pre-trained models to memorize and reproduce training data (Carlini et al., 2019, 2023; Nasr et al., 2023), presents significant challenges in terms of legal issues and sensitive data leakage. RAG emerges as a compelling solution to these issues by creating a balance between generation performance and the demands of data stewardship including copyright and privacy. Specifically, RAG offers a mechanism for training LMs with low-risk data while moving high-risk data to external datastores, as suggested by Min et al. (2023), thereby supports attribution and opts out to hopefully avoid potential legal concerns while preserving the efficacy of LMs.
@@
We start by building RIC-based RAG systems using popular open-sourced instruction-tuned LMs as generative models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and use newest Wikipedia articles (created later than November 1st, 2023) as datastore. Then adversarial prompts are developed to effectively extract nearly verbatim texts from the datastores of RAG models. We show that LMs with strong capabilities suffer from a high risk of disclosing context, and the vulnerability is exacerbated as the model size scales up from 7B to 70B. Moreover, our ablation studies reveal that instruction tuning makes LMs more prone to follow malicious instructions, and this can be evidence of failure in effectively utilizing contexts (Liu et al., 2024).
@@
2 Related Work
Retrieval-Augmented Generation. RAG (Lewis et al., 2020) has been widely studied in the NLG domain, such as kNN-LM (Khandelwal et al., 2019), DPR (Karpukhin et al., 2020), RALM (Guu et al., 2020), RETRO (Borgeaud et al., 2022) and REPLUG (Shi et al., 2023). We focus on a popular implementation of RAG - RIC-LM (Ram et al., 2023) that retrieves text chunks from a datastore and feeds them to an LM in context. There has been growing interest in analyzing data leakage problems of RAG systems, including customized GPTs. Huang et al. (2023) first conduct the study of privacy issues on kNN-LMs and show that incorporating private datastores leads to higher risks of data leakage from datastores. Yu et al. (2023) leverage prompt injection to cause file leakage of GPTs by asking them to download the uploaded files using ChatGPT’s code interpreter. Zeng et al. (2024) designed adversarial prompts to cause privacy leakage from external datastore, which is further studied in multi-turn interaction scenarios by Jiang et al. (2024b). Zyskind et al. (2023) propose secure multi-party computation that allows users to privately search a database. We are the first to comprehensively study data leakage problems on both open-sourced and production RAG systems and across multiple settings, and our attack on GPTs reached a 100% success rate without additional tools. Our work focuses on scenarios where datastores should be kept private, which can encompass an array of LM-integrated complex systems, e.g. distributing a customized non-parametric memory-based agent (Park et al., 2023; OpenAI, 2024) to third-party users (OpenAI, 2023); retrieving private yet high-quality data that the model creator does not desire to share with users (Brown et al., 2022); retrieving from pre-training corpora that are not well-sanitized so might contain personally identifiable information (PII) etc sensitive data (Elazar et al., 2023; Subramani et al., 2023).
@@ 4 Attacking Open-sourced LMs
RAG Setup.   We simulate a scenario where the service provider uses the latest Wikipedia content as the knowledge base. To construct the datastore, we collect 1165 recent Wikipedia English articles created after November 1st, 2023, with 1,569,000 words in total. We choose this recent cutoff date to ensure the models we used have not likely been trained on those Wikipedia texts so datastore content is out of the LMs’ knowledge. We use the Wikipedia API to automatically download the data and filter out articles less than 100 words. We use instruction-tuned LMs as the generative models. To test LMs across different sizes, we choose Llama2-Chat (7b/13b/70b) (Touvron et al., 2023), Vicuna (13b) (Chiang et al., 2023), Mistral-Instruct (7b) (Jiang et al., 2023) and Mixtral-Instruct (8x7b) (Jiang et al., 2024a), SOLAR (10.7b) (Kim et al., 2023), WizardLM (13b) (Xu et al., 2023a), Qwen1.5-Chat (72b) (Bai et al., 2023), and Platypus2-Instruct (70b) (Lee et al., 2023). Note that these LMs are not only instruction-tuned but also aligned with human preferences to ensure safety and helpfulness, thereby having the capability to detect and reject malicious prompts to some extent.
@@ 4.1 Ablation Studies
When do LMs tend to follow unexpected instructions? In practice, user queries are usually inserted at different positions of context windows as the conversation goes on, rather than our default setting where such queries are only appended at the end. Motivated by the shortcomings of LMs in effectively utilizing contexts (Liu et al., 2024; Wang et al., 2024; Anil et al., 2024), we hypothesize that LMs are more prone to follow instructions of context reconstruction that are near the beginning or end of the input context. We verify the hypothesis in two different settings: Adversarial prompt is inserted 1) at the beginning/end of the context window, and 2) in the middle of the context window. Note that it’s not a practical setting that’s adopted by current RAG systems, and the study simplifies the scenario and serves as a proof of concept.
@@
We set the number of retrieved chunks to 16 and the maximum length per chunk to 128, and inject the adversarial prompt to the 15 intervals between the chunks. From Figure 5 we observe a U-shaped curve where the reconstruction length reaches high values at both ends while reaching a minimum in the middle. Also, the minimum occurs at a position near the right end, indicating that a prompt injection near the beginning of the context tends to reconstruct more data from the context. We hypothesize that this phenomenon is an example of LMs struggling to process information in the middle of context as evidenced in (Wang et al., 2024). Many modern LMs, including our chosen Mistral-Instruct-7b, use RoPE (Su et al., 2024) for position encoding, which suffers from recency bias (Peysakhovich and Lerer, 2023), causing LMs to focus on the most recent information (end of context). Additionally, the causal attention mechanism in autoregressive LMs propagates information from left to right, making them focus more on distant information (beginning of context).
@@ 6 Conclusion
Limitations
As a proof of concept, we focus only on widely used Retrieval-In-Context RAG models with adversarial prompts, but leave efficient automated attack designs to other RAG implementations as future work. Moreover, we only propose an attack without a corresponding defense approach. Future work should consider designing an effective data protection method exploiting privacy-preserving training or inference without significant utility degradation of the RAG system.
Broader Impact

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) improves pre-trained models by incorporating external knowledge at test time to enable customized adaptation. We study the risk of datastore leakage in Retrieval-In-Context RAG Language Models (LMs). We show that an adversary can exploit LMs’ instruction-following capabilities to easily extract text data verbatim from the datastore of RAG systems built with instruction-tuned LMs via prompt injection. The vulnerability exists for a wide range of modern LMs that span Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and the exploitability exacerbates as the model size scales up. We also study multiple effects of RAG setup on the extractability of data, indicating that following unexpected instructions to regurgitate data can be an outcome of failure in effectively utilizing contexts for modern LMs, and further show that such vulnerability can be greatly mitigated by position bias elimination strategies. Extending our study to production RAG models GPTs, we design an attack that can cause datastore leakage with a 100% success rate on 25 randomly selected customized GPTs with at most 2 queries, and we extract text data verbatim at a rate of 41% from a book of 77,000 words and 3% from a corpus of 1,569,000 words by prompting the GPTs with only 100 queries generated by themselves. Code is available at this repository.
@@ 1 Introduction
In particular, the inherent propensity of large pre-trained models to memorize and reproduce training data (Carlini et al., 2019, 2023; Nasr et al., 2023), presents significant challenges in terms of legal issues and sensitive data leakage. The approach of RAG emerges as a compelling solution to these issues by creating a balance between generation performance and the demands of data stewardship including copyright and privacy. Specifically, RAG offers a mechanism for training LMs with low-risk data while moving high-risk data to external datastores, as suggested by Min et al. (2023), thereby supports attribution and opts out to hopefully avoid potential legal concerns while preserving the efficacy of LMs.
@@
We start by building RIC-based RAG systems using popular open-sourced instruction-tuned LMs as generative models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and use newest Wikipedia articles (created later than November 1st, 2023) as datastore. Then adversarial prompts are developed to effectively extract nearly verbatim texts from the datastores of RAG models. We show that LMs with strong capabilities suffer from a high risk of disclosing context, and the vulnerability is exacerbated as the model size scales up from 7B to 70B. Furthermore, our ablation studies indicate that instruction tuning increases the susceptibility of language models to follow malicious instructions. Our results also suggest such vulnerabilities might stem from the presence of position bias and a failure to effectively utilize contextual information (Liu et al., 2024). Motivated by these findings, we explore position-bias elimination strategies and propose that combining them with safety-aware prompts can effectively defend against prompt-injected data extraction attacks.
@@ 3 Attacking Open-sourced LMs
RAG Setup.   We simulate a scenario where the service provider uses the latest Wikipedia content as the knowledge base. To construct the datastore, we collect 1165 recent Wikipedia English articles created after November 1st, 2023, with 1,569,000 words in total. We choose this recent cutoff date to ensure the models we used have not likely been trained on those Wikipedia texts so datastore content is out of the LMs’ knowledge. We use the Wikipedia API to automatically download the data and filter out articles less than 100 words. We use instruction-tuned LMs as the generative models. To test LMs across different sizes, we choose Llama2-Chat (7b/13b/70b) (Touvron et al., 2023), Vicuna (13b) (Chiang et al., 2023), Mistral-Instruct (7b) (Jiang et al., 2023) and Mixtral-Instruct (8x7b) (Jiang et al., 2024), SOLAR (10.7b) (Kim et al., 2023), WizardLM (13b) (Xu et al., 2023a), Qwen1.5-Chat (72b) (Bai et al., 2023), and Platypus2-Instruct (70b) (Lee et al., 2023). Note that these LMs are not only instruction-tuned but also aligned with human preferences to ensure safety and helpfulness, thereby having the capability to detect and reject malicious prompts to some extent.
@@ 3.1 Ablation Studies
When do LMs tend to follow unexpected instructions? In practice, user queries are usually inserted at different positions of context windows as the conversation goes on, rather than our default setting where such queries are only appended at the end. Motivated by the shortcomings of LMs in effectively utilizing contexts (Liu et al., 2024; Wang et al., 2024a; Anil et al., 2024), we hypothesize that LMs are more prone to follow instructions of context reconstruction that are near the beginning or end of the input context. We verify the hypothesis in two different settings: Adversarial prompt is inserted 1) at the beginning/end of the context window, and 2) in the middle of the context window. Note that it’s not a practical setting that’s adopted by current RAG systems, and the study simplifies the scenario and serves as a proof of concept.
@@
We set the number of retrieved chunks to 16 and the maximum length per chunk to 128, and inject the adversarial prompt to the 15 intervals between the chunks. From Figure 5 we observe a U-shaped curve where the reconstruction length reaches high values at both ends while reaching a minimum in the middle. Also, the minimum occurs at a position near the right end, indicating that a prompt injection near the beginning of the context tends to reconstruct more data from the context. We hypothesize that this phenomenon is an example of LMs struggling to process information in the middle of context as evidenced in (Wang et al., 2024a). Many modern LMs, including our chosen Mistral-Instruct-7b, use RoPE (Su et al., 2024) for position encoding, which suffers from recency bias (Peysakhovich and Lerer, 2023), causing LMs to focus on the most recent information (end of context). Additionally, the causal attention mechanism in autoregressive LMs propagates information from left to right, making them focus more on distant information (beginning of context).
3.2 Mitigation Strategies
In response to the prompt-injected data extraction attacks discussed previously, we investigate two mitigation strategies aimed at addressing the the vulnerabilities. These strategies are designed to reduce the model’s susceptibility to prompt injection by enhancing its ability to distinguish between legitimate and adversarial prompts.
We conducted experiments using the Llama3 8b Instruct model, replicating the procedures detailed in Section 3. The experimental setup adheres to the configurations specified in the subsections “RAG Setup” and “Attack Setup”. For evaluation, we employed the Rouge-L and BERTScore metrics, and additionally included the Reconstruction Rate ($R$) that measures the effectiveness of the extracted chunks in reconstructing the original text data. It is calculated as the ratio of the total length of the concatenated, deduplicated text chunks to the length of the original text data. Formally, let:
  • $O$ denote the original text data in the datastore.
  • $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ represent the set of extracted chunks.
  • $\mathcal{C}^{\prime}=\{c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{m}\}$ denote the deduplicated set of chunks obtained from $\mathcal{C}$.
  • $|X|$ denote the length of text $X$.
The Reconstruction Rate ($R$) is then defined as: $R=\frac{\sum_{i=1}^{m}|c^{\prime}_{i}|}{|O|}$. A higher Reconstruction Rate indicates that a larger portion of the original text has been successfully reconstructed from the extracted chunks, whereas a lower rate suggests less effective reconstruction.
We compare the baseline RAG configuration—without any mitigation methods—to two proposed mitigation strategies: Safety-Aware Prompt and Position Bias Elimination.
Method Rouge-L BERTScore Reconstruction Rate
Baseline 91.29 97.08 88.88
Safety-aware Prompt 91.13 97.13 87.57
PINE 71.29 92.14 58.03
Safety-aware Prompt + PINE 67.25 91.26 52.34
Table 3: Comparison of mitigation strategies against prompt-injected data extraction attacks. Lower scores indicate better mitigation performance.
3.2.1 Safety-Aware Prompt
Our baseline mitigation strategy involves incorporating a safety-aware system prompt that explicitly instructs the model to refrain from reproducing content from the retrieved context. Specifically, we employ the following system prompt: “Do not repeat any content from the context.” As shown in Table 3, this approach results in a slight decrease in the reconstruction rate. However, the overall reduction in vulnerability to prompt-injected data extraction attacks is marginal, indicating that this strategy alone offers limited protection.
3.2.2 Position Bias Elimination
Our prior observations highlight the impact of position bias on data reconstruction, leading us to hypothesize that the model’s vulnerability arises from its inability to distinguish malicious instructions from the system prompt and legitimate retrieved documents. To address this issue, we implement position bias elimination strategies, specifically utilizing Position-Insensitive Encoding (PINE) (Wang et al., 2024b) as a representative technique. PINE enables the explicit grouping of text segments, allowing the language model to process all segments within a group equally while distinguishing them from those outside the group. In our defense mechanism, we apply this approach by grouping the user query and the retrieved documents together, thereby isolating them from the system prompt. The input is restructured as [system prompt, [retrieved doc 1, retrieved doc 2, user query], <EOS>], ensuring that the retrieved documents and user query are attended to equally while the system prompt remains separate. This separation reduces the likelihood of the model inadvertently following adversarial instructions embedded within the prompt. The results in Table 3 demonstrate that PINE significantly lowers the reconstruction rates, confirming its effectiveness as a standalone mitigation strategy.
3.2.3 Combined Strategy: Safety-Aware Prompt and PINE
Building on the individual strengths of the safety-aware prompt and PINE, we evaluate the combined application of both mitigation methods. This integrated approach yields the most significant improvements, achieving the lowest reconstruction rates across all evaluated metrics, including Rouge-L and BERTScore. The combined strategy effectively addresses both the instruction-following behavior of LLMs and the position bias within the context window. The robustness gain is empirically justifiable, as the attention scores are recalculated using bi-directional attention, and then sorted to ensure the model treats all instructions equally. This process prevents the model from disproportionately following the final instruction, maintaining position-invariance during inference.
@@ 4 Attacking Production LMs
5 Related Work
Retrieval-Augmented Generation. RAG (Lewis et al., 2020) has been widely studied in the NLG domain, such as kNN-LM (Khandelwal et al., 2019), DPR (Karpukhin et al., 2020), RALM (Guu et al., 2020), RETRO (Borgeaud et al., 2022) and REPLUG (Shi et al., 2023). We focus on a popular implementation of RAG - RIC-LM (Ram et al., 2023) that retrieves text chunks from a datastore and feeds them to an LM in context. There has been growing interest in analyzing data leakage problems of RAG systems, including customized GPTs. Huang et al. (2023) first conduct the study of privacy issues on kNN-LMs and show that incorporating private datastores leads to higher risks of data leakage from datastores. Yu et al. (2023) leverage prompt injection to cause file leakage of GPTs by asking them to download the uploaded files using ChatGPT’s code interpreter, while our proposed attack on GPTs reached a 100% success rate without additional tools. Zyskind et al. (2023) propose secure multi-party computation that allows users to privately search a database.
The most related study to our work is conducted by Zeng et al. (2024), who designed adversarial prompts to cause privacy leakage from external datastore. However, Zeng et al. (2024) did not perform experiments on production-level RAG systems, thereby limiting the practical implications. Secondly, although they demonstrate the potential for extracting private data from open-source RAG systems, their investigation does not extend to analyzing the underlying reasons or the impact of various RAG configurations–such as model size, the position of query in context window, and the distinction between seen and unseen data–on data leakage. In contrast, we comprehensively study data leakage problems on both open-sourced and production RAG systems and across multiple settings, leading to effective mitigation strategies and providing a more comprehensive understanding of how different RAG settings influence data leakage vulnerabilities.
Our work focuses on scenarios where datastores should be kept private, which can encompass an array of LM-integrated complex systems, e.g. distributing a customized non-parametric memory-based agent (Park et al., 2023; OpenAI, 2024) to third-party users (OpenAI, 2023); retrieving private yet high-quality data that the model creator does not desire to share with users (Brown et al., 2022); retrieving from pre-training corpora that are not well-sanitized so might contain personally identifiable information (PII) etc sensitive data (Elazar et al., 2023; Subramani et al., 2023).
@@ 6 Conclusion
Ethics Considerations